# Unlocking massively parallel spectral proper orthogonal decompositions in the PySPOD package

### v1

@@ 3.2.2 PySPOD I/O
In PySPOD, we take advantage of the N–N model’s good performance, but like authors of Yu et al. (2012), slightly modify it to be N–M (generally M $<$ N) mapping, where M is not to be confused with ‘italic’ $M$ adopted in section 2.3. Such a setting, where N is the number of processes and M is the number of files, is better suited to our application characteristics. We use a simple two-phase I/O remnant of that proposed in Del Rosario et al. (1993) back in 1993.
@@
    PySPOD is designed to analyze large datasets, which are often split over multiple files. One such example is the Climate Data Store (CDS), which provides rich climate datasets Hersbach et al. (2018), and that has been used in the scalability analyses presented in sections 5.2 and 5.3. CDS, however, limits the amount of data that can be downloaded in a single request. The resulting datasets’ size can be in the order of tens of terabytes split over several hundred files. Another example arises from computational fluid dynamics (CFD) simulators such as SSDC, Nektar++ and Charles Parsani et al. (2021); Moxey et al. (2020); Brès et al. (2017). Such software often writes output using one file per requested timestep. Considering the data size and PySPOD’s competitive scalability performance, we decided against converting the data to a common format (such as HDF5), as the conversion could be more expensive than the analysis, and we implemented several native readers for different data formats.
@@
    Using a two-phase I/O results in fewer and more contiguous requests to the storage system, which is preferred on a parallel file system. For illustration, most often, the data is split over multiple files, where each file represents a different range of timesteps, and the spatial coordinates follow the first dimension. In our design, each process reads contiguous data from at least one input file, i.e., all spatial dimensions for a subset of timesteps (or, more generally, the first dimension, i.e., time), which is good for performance. An alternative design, such as using MPI-IO to immediately read only a subset of spatial variables for all timesteps, may result in every process accessing every file, depending on the MPI implementation used. In fact, such a reader is also implemented in PySPOD; however, based on poor performance results obtained in early experimentation, we decided to extend and optimize the two-phase reader.
@@
    Separating I/O from MPI removes the risk of performance degradation due to an underoptimized MPI distribution. As discussed in section 3.2.1, MPI I/O’s performance may depend heavily on the MPI distribution used and its knowledge of the underlying I/O system. By separating I/O and data distribution, we read the data from the disk efficiently and then explicitly distribute it with MPI. Since we use point-to-point communication, which is at the very core of MPI, it is much less likely that those functions will perform poorly.
@@
To address the memory capacity limitations of PySPOD, we implemented several precautions. One such measure was to read data in chunks, with each process reading approximately 256 MB of data (value determined empirically). This helps reducing memory overhead since processing each chunk requires twice the memory, as time-split data being sent can only be deallocated once it is received by the target process, and space-split data is being received simultaneously.
@@
Additionally, we implemented reading in such a way that one process first seeks through the metadata to identify which range of timesteps is available in which file. This information, broadcast with MPI, significantly reduces the overhead associated with N processes opening all M files.
@@ 4 Datasets
The datasets adopted to test the parallel SPOD algorithm consists of fluid mechanics, and geophysical data. The former uses jet data produced by high-fidelity large-eddy simulation (LES), and is described in section 4.1. The latter uses fifth-generation reanalysis data (ERA5) produced by ECMWF Hersbach et al. (2020), and is described in section 4.2, along with associated SPOD results.
@@ 4.1 SPOD analysis of fluid mechanics datasets
In our first example, we analyze the data generated by Yeung et al. Yeung et al. (2022) using the solver Charles Brès et al. (2017), from LES of a supersonic twin-rectangular jet at a Reynolds number of $Re\approx 10^{6}$ based on the jet exit conditions and the equivalent nozzle diameter. The simulation was previously validated by Brès et al. Brès et al. (2021) against the companion experiments of Samimy et al. Samimy et al. (2023). The time-resolved data consist of 20,000 snapshots of the 3D flow field, interpolated onto a Cartesian grid and saved at a time interval of $\Delta t=0.2$. Each snapshot records five primitive variables: density ($\rho$), velocities ($u,v,w$), and temperature ($T$), in single precision, for a total of $N_{x}\times N_{y}\times N_{z}\times N_{\mathrm{var}}=625\times 270\times 344\times 5$ data points per snapshot. Storage of the database requires 18,392 GB on disk in HDF5 format, or 43,251 GB once loaded into memory.
@@
such that the inner product $\expectationvalue{\mathbf{q}_{1},\mathbf{q}_{2}}=\mathbf{q}_{1}^{\mathrm{H}}\mathbf{W}\mathbf{q}_{2}$ induces the compressible energy norm Chu (1965). Here, $\gamma=1.4$ is the adiabatic index; $M_{j}=1.5$ is the jet Mach number. We select a block size of $N_{f}=256$, with 50% overlap, giving $L=155$ blocks.
The premultiplied SPOD eigenvalue spectra are reported in figure 3(a). The leading eigenvalues show a prominent peak at $f\approx 0.2$, where $f$ is nondimensionalized by the nozzle height and the ambient speed of sound. This peak corresponds to the fundamental screech tone. Screeching of the supersonic twin-rectangular jet stems from acoustic resonance between the nozzle and the shock cells, and has been observed both experimentally Jeun et al. (2022); Samimy et al. (2023) and numerically Brès et al. (2021, 2022); Yeung et al. (2022) to occur at a similar frequency. The peak at $f\approx 0.2$ persists at least to the second eigenvalue. Large separations between the first, second, and third eigenvalues—termed low-rank behavior Schmidt et al. (2018)—in the frequency range $0.2\lesssim f\lesssim 0.3$ signal the presence of coherent structures arising from underlying physical instabilities, in this case the screech mechanism.
@@
Isovalues of $\mathrm{Re}\{{\bf\it\phi}_{p}\}=\pm 0.0005$ are chosen to highlight the 3D structure of the far-field acoustic waves. Each row in figure 4 shows one mode. The left and right columns, 4(a,c,e,g) and 4(b,d,f,h), also display the cross-sectional views of the planes $z=1.3$ and $y=-0.25$, respectively. These provide insights into the wavepackets within the jet plume, which are well-known to be efficient sources of noise Jordan and Colonius (2013). In figure 4(a,b), mode 1 recovers near-perfect antisymmetry about the major-axis plane, $y=0$, and symmetry about the minor-axis plane, $z=0$. In contrast, mode 2 in 4(c,d) is clearly antisymmetric about both planes. Mode 4 in 4(g,h), on the other hand, is symmetric about both planes. The symmetry of mode 3 is difficult to ascertain visually from 4(e,f). In particular, the strong in-phase (4(b,h)) and out-of-phase (4(d)) coupling between the twin jets observed in modes 1, 2, and 4 appear to be lost in mode 3 (4(f)). This is due to insufficient statistical convergence.
The twin-rectangular nozzle configuration is geometrically invariant with respect to reflection about the major and minor axes. It thus belongs in the dihedral group $D_{2}$. Such geometrical symmetries are broken by turbulence; however, as figures. 2(b) and 4 demonstrate, they can reappear in the statistics. To investigate the extent to which the leading SPOD modes at all frequencies exhibit symmetry or antisymmetry, we perform $D_{2}$-symmetry decomposition Yeung et al. (2022) of each mode into four components,
${\bf\it\phi}={\bf\it\phi}_{\mathrm{SS}}+{\bf\it\phi}_{\mathrm{SA}}+{\bf\it\phi}_{\mathrm{AS}}+{\bf\it\phi}_{\mathrm{AA}}\,,$ (21)
where the subscripts denote symmetry (S) or antisymmetry (A) about the major (first letter) and minor (second letter) axes Rodríguez et al. (2018). The importance of each symmetry can be quantified by projecting the mode onto its symmetry components. We define the normalized projection coefficient,
$\alpha=\absolutevalue{\frac{\expectationvalue{{\bf\it\phi},{\bf\it\phi}_{\mathrm{S/A}}}}{\expectationvalue{{\bf\it\phi},{\bf\it\phi}}}}\,.$ | . (22)
Intuitively, $\alpha$ measures the relative energy of each component as defined by the compressible energy norm in equation (19).
In figure 3(b), we color the first four eigenvalues at each frequency by the symmetry component that contains the maximum relative energy. The percent opacity is set to $\alpha$; in other words, 100% opacity indicates the SPOD mode perfectly recovers a particular symmetry. At $f=0.21$, the colors indicate that modes one through four are dominated by the AS, AA, SA, and SS symmetries, respectively. These symmetries confirm our visual observations of the modes in figure 4. The dominant symmetry can vary between mode numbers, but also across frequencies. For instance, the leading mode switches from the AS symmetry at $f\approx 0.2$ to the SA symmetry in the range $0.03\lesssim f\lesssim 0.2$. The AS symmetry then re-emerges below $f\approx 0.03$. Variations in dominant symmetries reflect the property that, in general, SPOD modes at distinct frequencies represent distinct physical structures Schmidt et al. (2018). The degree of symmetry recovery, $\alpha$, falls off with increasing mode number as a consequence of incomplete convergence. On the other hand, the decay of $\alpha$ with frequency in the range $f\gtrsim 0.2$ suggests that higher-frequency modes, which have more limited spatial support, are less likely to establish shear-layer coupling. Lack of coupling in turn prevents modes from achieving symmetry.
A comprehensive discussion of twin-rectangular jet dynamics exceeds the scope of this work. However, our findings illustrate some of the physical insights that can only be gleaned from a 3D modal analysis, which PySPOD vastly accelerates.
Figure 3: Premultiplied SPOD eigenvalue spectra of the twin-rectangular jet. The spectra in (a) fade from black to white with increasing mode number. Only the first four eigenvalues at each frequency are displayed in (b), where they are colored by the dominant $D_{2}$-symmetry component of the respective SPOD modes, with higher opacity denoting greater symmetry recovery, $\alpha$. Modes corresponding to the highlighted ($\bullet$) eigenvalues at $f=0.21$ are reported in figure 4.
@@ 4.2 SPOD analysis of geophysical datasets
In our second example, we use data from ERA5, a fifth-generation reanalysis dataset produced by the ECMWF that combines model data with global observations using the laws of physics to create a complete and consistent global climate and weather dataset. We obtained this data from the Climate Data Store (CDS) Hersbach et al. (2018).
In particular, we consider the horizontal speed of air moving towards the east on 37 pressure levels (i.e., vertical levels) spanning from January 1940 to December 2022 Hersbach et al. (2018). This quantity is also referred to as U component of the horizontal wind velocity, and its unit is meters per second. The dataset contains 727,584 time snapshots on a 3D grid of dimension 1440$\times$721$\times$37. The total size of this dataset is 51,745 GB, corresponding to 27.9 trillion data points and 103,490 GB in memory when using single-precision floating-point, as for this example, and 199,452 GB in double-precision. In figure 5, we depict the horizontal wind velocity adopted for pressure level 1 (top left), 12 (top right), 24 (bottom left), and 37 (bottom right). dataset.
@@
The analysis of this dataset aims to capture the quasi-biennial oscillation (QBO) that has an approximate period of 2 to 2.5 years, as reported in Schmidt et al. (2019). This atmospheric oscillation is characterized by quasi-periodic reversals of the zonal-mean zonal winds in the equatorial stratosphere – see also Baldwin et al. (2001). QBO has important implications on teleconnections, influencing weather patterns in the Northern Hemisphere, and the tropics (including tropical precipitation) Gray et al. (2018). Its most distinctive feature is a latitudinal band of the U component of the wind velocity in the tropical region (±$\pm 20^{\circ}$ latitude).
Indeed, figure 6 shows the eigenvalue spectra for the U component of the wind velocity, where the leading eigenvalue show a prominent peak at period $T=912.5$ days. This peak corresponds to the QBO, and shows as this phenomenon exhibits low-rank behaviour, since the energy of the leading eigenvalue is remarkably separated from the other eigenvalues. Figure 7 shows the highly-coherent three-dimensional structure of the leading mode. This manifests as a latitudinal band of the U component of the wind velocity in the tropical region, that is the signature of QBO. The results are consistent with those in Schmidt et al. (2019), albeit the data adopted there was ERA-20C Poli et al. (2016), that has significantly coarser spatial and temporal resolution, but longer coverage (from 1900 to 2010). In our results, we also observe several high-frequency peaks in the eigenvalues (periods shorter than 1 day) thanks to the improved temporal resolution of our dataset.
@@ 5.1 Scalability setup
To test the scalability of the PySPOD library, we used the Shaheen II supercomputer, a Cray XC40 system hosted by King Abdullah University of Science and Technology (KAUST) Hadri et al. (2015). Each of Shaheen’s 6,174 nodes features two Intel Haswell (Xeon E5-2698v3) CPUs with 16 cores each and 128 GB of memory. The nodes are connected via a Cray Aries interconnect with Dragonfly topology. While the nodes were allocated in exclusive mode by the scheduler, the network was shared with other users, as is typical in a production environment. To alleviate the memory capacity bottleneck and maximize the I/O bandwidth, we run PySPOD with 4 processes per node.
For storage, we utilized the Cray Sonexion® 2000 Storage System, which offers over 16 PB of usable capacity using 72 Scalable Storage Units (SSUs), 144 Object Storage Services (OSSs) and 144 Object Storage Targets (OSTs). The system features 5,988 4 TB disks. The theoretical performance of this storage exceeds 500 GB/s Hadri et al. (2017), and published application benchmarks have shown that applications such as WRF can achieve a bandwidth of 25–35 GB/s, including communication, or 65 GB/s of raw I/O performance measured on aggregators when using 144 OSTs and 4 MPI ranks per node. Similar results have been obtained with NPB benchmarks Markomanolis et al. (2017).
@@
We report the following timings for PySPOD: time for I/O (combined reading and writing of 5 leading modes to disk), computation of the DFT (step 2 in figure 1), computation of the inner product and associated matrix $\hat{\mathbf{M}}$ (step 3 in figure 1), eigenvalue decomposition (step 5 in figure 1, and computation of the SPOD modes (just prior to step 6 in figure 1. Given the variability of performance on Dragonfly networks Chunduri et al. (2017) and shared file systems Lofstead et al. (2010), we report arithmetic averages based on 5 repetitions.
The main development branch of PySPOD (commit TODO:6) was utilized, along with the following packages: mpi4py 3.1.4, netCDF4 1.6.3, NumPy 1.24.2, SciPy 1.10.1, and Xarray 2023.2.0. Cray-provided modules for Python 3.10.1 and Cray MPICH 7.7.18 were used.

### v2

@@ 3.2.2 PySPOD I/O
In PySPOD, we take advantage of the N–N model’s good performance, but like authors of [37], slightly modify it to be N–M (generally M $<$ N) mapping, where M is not to be confused with ‘italic’ $M$ adopted in section 2.3. Such a setting, where N is the number of processes and M is the number of files, is better suited to our application characteristics. We use a simple two-phase I/O remnant of that proposed in [21] back in 1993.
@@
    PySPOD is designed to analyze large datasets, which are often split over multiple files. One such example is the Climate Data Store (CDS), which provides rich climate datasets [38], and that has been used in the scalability analyses presented in sections 5.2 and 5.3. CDS, however, limits the amount of data that can be downloaded in a single request. The resulting datasets’ size can be in the order of tens of terabytes split over several hundred files. Another example arises from computational fluid dynamics (CFD) simulators such as SSDC, Nektar++ and Charles [39, 40, 41]. Such software often writes output using one file per requested timestep. Considering the data size and PySPOD’s competitive scalability performance, we decided against converting the data to a common format (such as HDF5), as the conversion could be more expensive than the analysis, and we implemented several native readers for different data formats.
@@
    Using a two-phase I/O results in fewer and more contiguous requests to the storage system, which is preferred on a parallel file system. For illustration, most often, the data is split over multiple files, where each file represents a different range of timesteps, and the spatial coordinates follow the first dimension. In our design, each process reads contiguous data from at least one input file, i.e., all spatial dimensions for a subset of timesteps (or, more generally, the first dimension, i.e., time), which is good for performance. An alternative design, such as using MPI-IO to immediately read only a subset of spatial variables for all timesteps, may result in every process accessing every file, depending on the MPI implementation used. As we discussed in section 3.2.1, MPI-IO has been shown to perform poorly on Lustre file systems [30]. In fact, an MPI-IO reader is also implemented in PySPOD; however, we also obtained poor performance results in early experimentation; hence, we decided to extend and optimize the two-phase reader.
@@
    Separating I/O from MPI removes the risk of performance degradation due to an underoptimized MPI distribution. As discussed in section 3.2.1, MPI-IO’s performance may depend heavily on the MPI distribution used and its knowledge of the underlying I/O system. By separating I/O and data distribution, we read the data from the disk efficiently and then explicitly distribute it with MPI. Since we use point-to-point communication, which is at the very core of MPI, it is much less likely that those functions will perform poorly.
@@
To address the memory capacity limitations of PySPOD, we implemented several precautions. One such measure was to read data in chunks, with each process reading approximately 256 MB of data (value determined empirically). This helps reducing memory overhead since processing each chunk requires twice the memory, as time-split data being sent can only be deallocated once it is received by the target process, and space-split data is being received simultaneously. The chunk size should be as small as possible to reduce the memory overhead of using auxiliary memory but large enough to hide the I/O latency. On our system, 256 MB chunks correspond to roughly 6% of total memory per CPU core.
@@
Moreover, we implemented reading in such a way that one process first seeks through the metadata to identify which range of timesteps is available in which file. This information, broadcast with MPI, significantly reduces the overhead associated with N processes opening all M files.
@@ 4 Datasets
The datasets adopted to test the parallel SPOD algorithm consists of fluid mechanics, and geophysical data. The former uses jet data produced by high-fidelity large-eddy simulation (LES), and is described in section 4.1. The latter uses fifth-generation reanalysis data (ERA5) produced by ECMWF [42], and is described in section 4.2, along with associated SPOD results.
@@ 4.1 SPOD analysis of fluid mechanics datasets
In our first example, we analyze the data generated by Yeung et al. [43] using the solver Charles [41], from LES of a supersonic twin-rectangular jet at a Reynolds number of $Re\approx 10^{6}$ based on the jet exit conditions and the equivalent nozzle diameter. The simulation was previously validated by Brès et al. [44] against the companion experiments of Samimy et al. [45]. The time-resolved data consist of 20,000 snapshots of the 3D flow field, interpolated onto a Cartesian grid and saved at a time interval of $\Delta t=0.2$. Each snapshot records five primitive variables: density ($\rho$), velocities ($u,v,w$), and temperature ($T$), in single precision, for a total of $N_{x}\times N_{y}\times N_{z}\times N_{\mathrm{var}}=625\times 270\times 344\times 5$ data points per snapshot. Storage of the database requires 18,392 GB on disk in HDF5 format, or 23,220 GB once loaded into memory.
@@
such that the inner product $\expectationvalue{\mathbf{q}_{1},\mathbf{q}_{2}}=\mathbf{q}_{1}^{\mathrm{H}}\mathbf{W}\mathbf{q}_{2}$ induces the compressible energy norm [46]. Here, $\gamma=1.4$ is the adiabatic index; $M_{j}=1.5$ is the jet Mach number. We select a block size of $N_{f}=256$, with 50% overlap, giving $L=155$ blocks.
The premultiplied SPOD eigenvalue spectra are reported in figure 3. The leading eigenvalues show a prominent peak at $f\approx 0.2$, where $f$ is nondimensionalized by the nozzle height and the ambient speed of sound. This peak corresponds to the fundamental screech tone. Screeching of the supersonic twin-rectangular jet stems from acoustic resonance between the nozzle and the shock cells, and has been observed both experimentally [47, 45] and numerically [44, 48, 43] to occur at a similar frequency. The peak at $f\approx 0.2$ persists at least to the second eigenvalue. Large separations between the first, second, and third eigenvalues—termed low-rank behavior [49]—in the frequency range $0.2\lesssim f\lesssim 0.3$ signal the presence of coherent structures arising from underlying physical instabilities, in this case the screech mechanism.
@@
Isovalues of $\mathrm{Re}\{{\bf\it\phi}_{p}\}=\pm 0.0005$ are chosen to highlight the 3D structure of the far-field acoustic waves. Each row in figure 4 shows one mode. The left and right columns, 4(a,c,e,g) and 4(b,d,f,h), also display the cross-sectional views of the planes $z=1.3$ and $y=-0.25$, respectively. These provide insights into the wavepackets within the jet plume, which are well-known to be efficient sources of noise [50]. In figure 4(a,b), mode 1 recovers near-perfect antisymmetry about the major-axis plane, $y=0$, and symmetry about the minor-axis plane, $z=0$. In contrast, mode 2 in 4(c,d) is clearly antisymmetric about both planes. Mode 4 in 4(g,h), on the other hand, is symmetric about both planes. The symmetry of mode 3 is difficult to ascertain visually from 4(e,f). In particular, the strong in-phase (4(b,h)) and out-of-phase (4(d)) coupling between the twin jets observed in modes 1, 2, and 4 appear to be lost in mode 3 (4(f)). This is due to insufficient statistical convergence.
A comprehensive discussion of twin-rectangular jet dynamics exceeds the scope of this work. However, our findings illustrate some of the physical insights that can be gleaned from a 3D modal analysis, which PySPOD vastly accelerates.
Figure 3: Premultiplied SPOD eigenvalue spectra of the twin-rectangular jet. The spectra fade from black to white with increasing mode number. Modes corresponding to the highlighted ($\bullet$) eigenvalues at $f=0.21$ are reported in figure 4.
@@ 4.2 SPOD analysis of geophysical datasets
In our second example, we use data from ERA5, a fifth-generation reanalysis dataset produced by the ECMWF that combines model data with global observations using the laws of physics to create a complete and consistent global climate and weather dataset. We obtained this data from the Climate Data Store (CDS) [38].
In particular, we consider the horizontal speed of air moving towards the east on 37 pressure levels (i.e., vertical levels) spanning from January 1940 to December 2022 [38]. This quantity is also referred to as U component of the horizontal wind velocity, and its unit is meters per second. The dataset contains 727,584 time snapshots on a 3D grid of dimension 1440$\times$721$\times$37. The total size of this dataset is 51,745 GB, corresponding to 27.9 trillion data points and 103,490 GB in memory when using single-precision floating-point, as for this example, and 199,452 GB in double-precision. In figure 5, we depict the horizontal wind velocity adopted for pressure level 1 (top left), 12 (top right), 24 (bottom left), and 37 (bottom right). dataset.
@@
This SPOD configuration was chosen to capture low-frequency atmospheric modes. More specifically, the analysis of this dataset aims to capture the quasi-biennial oscillation (QBO) that has an approximate period of 2 to 2.5 years, as reported in [6]. This atmospheric oscillation is characterized by quasi-periodic reversals of the zonal-mean zonal winds in the equatorial stratosphere – see also [51]. QBO has important implications on teleconnections, influencing weather patterns in the Northern Hemisphere, and the tropics (including tropical precipitation) [52]. Its most distinctive feature is a latitudinal band of the U component of the wind velocity in the tropical region (±$\pm 20^{\circ}$ latitude).
Indeed, figure 6 shows the eigenvalue spectra for the U component of the wind velocity, where the leading eigenvalue show a prominent peak at period $T=912.5$ days. This peak corresponds to the QBO, and shows as this phenomenon exhibits low-rank behaviour, since the energy of the leading eigenvalue is remarkably separated from the other eigenvalues. Figure 7 shows the highly-coherent three-dimensional structure of the leading mode. This manifests as a latitudinal band of the U component of the wind velocity in the tropical region, that is the signature of QBO. The results are consistent with those in [6], albeit the data adopted there was ERA-20C [53], that has significantly coarser spatial and temporal resolution, but longer coverage (from 1900 to 2010). In our results, we also observe several high-frequency peaks in the eigenvalues (periods shorter than 1 day) thanks to the improved temporal resolution of our dataset.
@@ 5.1 Scalability setup
To test the scalability of the PySPOD library, we used the Shaheen II supercomputer, a Cray XC40 system hosted by King Abdullah University of Science and Technology (KAUST) [54]. Each of Shaheen’s 6,174 nodes features two Intel Haswell (Xeon E5-2698v3) CPUs with 16 cores each and 128 GB of memory. The nodes are connected via a Cray Aries interconnect with Dragonfly topology. While the nodes were allocated in exclusive mode by the scheduler, the network was shared with other users, as is typical in a production environment. To alleviate the memory capacity bottleneck and maximize the I/O bandwidth, we run PySPOD with 4 MPI processes per compute node. We do not use multithreading.
For storage, we utilized the Cray Sonexion® 2000 Storage System, which offers over 16 PB of usable capacity using 72 Scalable Storage Units (SSUs), 144 Object Storage Services (OSSs) and 144 Object Storage Targets (OSTs). The system features 5,988 4 TB disks. The theoretical performance of this storage exceeds 500 GB/s [55], and published application benchmarks have shown that applications such as WRF can achieve a bandwidth of 25–35 GB/s, including communication, or 65 GB/s of raw I/O performance measured on aggregators when using 144 OSTs and 4 MPI ranks per node. Similar results have been obtained with NPB benchmarks [56].
@@
We report the following timings for PySPOD: time for I/O (combined reading and writing of 5 leading modes to disk), computation of the DFT (step 2 in figure 1), computation of the inner product and associated matrix $\hat{\mathbf{M}}$ (step 3 in figure 1), eigenvalue decomposition (step 5 in figure 1, and computation of the SPOD modes (just prior to step 6 in figure 1. Given the variability of performance on Dragonfly networks [57] and shared file systems [58], we report arithmetic averages based on 5 repetitions.
We employed PySPOD release 2.0.0, along with the following packages: mpi4py 3.1.4, netCDF4 1.6.3, NumPy 1.24.2, SciPy 1.10.1, and Xarray 2023.2.0. Cray-provided modules for Python 3.10.1 and Cray MPICH 7.7.18.